# A projection method for particle resampling

### v2

@@ 4.2 Particle grids and finite element order
The phase space continuum grids in this work are on regular 90 degree lattice with an option for simple r-refinement in velocity space. The particle grids are defined with a Cartesian particle grid, of a fixed size, in each phase space cell, similar to Lapenta (Figure 1, [25]). For simplicity, the original grid is used for resampling in this work and adaptivity strategies are left for future work. As discussed in §3.1, with periodic boundary conditions in the spatial dimension and natural boundary conditions in velocity dimension we use at least $Q+1$ particles in each dimension in each phase space cell for stability of the pseudoinverse, but we found that using more particles per cell is often desirable for noise reduction.
@@ 5 Finite element version of the direct remap method
An approach similar to our method that aligns low-order continuum space points with the particle points allows for a direct map from the grid to particles. A cell-centered finite difference version of this approach was developed by Colella et al. on Cartesian grids [41, 30], and with phase-space adaptive mesh refinement (AMR) [31]. These experiments are intended to observe the qualitative effect of high-order grid to particle mapping. Quantitatively, the direct map approach conserve energy to only four digits while the energy is conserved to all double precision digits with $Q2$ in velocity (see data the src/A.01 directory for $Q1$ and in src/A.0001 for $Q2$ in the data repository). The pseudoinverse projection also allows, and in fact requires, more that one particle per phase-space cell, which is common in practice and allows for adjusting the smoothing of the distribution that reduces weight variance in a spatial cell, which is a common criterial for resampling algorithms [10].
Our test harness supports a direct map method in a vertex-centered, finite element context. Linear $Q1$ elements are used in the velocity and space dimension on Cartesian grids. Particles are placed at vertices and the domain is doubly periodic, which is immaterial given that there should be negligible density at the velocity boundary. The salient feature of this construct is that the particle mass matrix is the identity, $\overline{M}_{p}=I$, and the pseudoinverse vanishes.
Direct map resampling is tested with a linear Landau damping, $A=0.01$, problem with a $128\times 256$ particle grid and $V_{max}=6.0$ and results in Figure 2 agree well with amplitude reported in [30] (Figure 3.1).
Figure 2: Particle electric field amplitude with linear Landau damping, $A=0.01$, with no resampling, a direct remapping finite element version of Myers et al., and the projection method
$Q_{1}-Q_{1}$ is used for the projection method with one particle per cell, or vertex, to mimic the direct map method, and the projection method uses a $2\times 2$ particle grid per cell and half as many cells in each dimension to maintain the same number of particles for all three tests and to ensure stability of the pseudo-inverse. This test show a modest improvement in the behavior of the solver, which is simply an observation of the effect of the high-order mapping to back to particles (linear instead of the constant of the direct map). This data also shows the efficacy of resampling in that it suppresses the noise observed without resampling for the $A=0.01$ case.
@@ 6 Numerical experiments with pseudoinverse resampling
This section investigates results of the new projection resampling method on a linear and a nonlinear Landau damping tests. An additional algorithmic feature used in these experiments is to use an ad-hoc method to enforce continuous, $C^{0}$ electric field. One can use an $H(div)$ Poisson solver and this is a subject of future work, but in this work the electric field computed at particles, with a $C^{0}$ potential field, is not $C^{0}$, which can cause particles to artificially reflect off of cell boundaries and with the highly regular grids used here large instabilities result with $Q2$ elements. These instabolites are at least partialy due to lack of a quite start method in this work [36].
To fix this problem the electric field is projected to the vertices and then back to the particles. This method precludes energy conservation, but there are other sources of loss of energy conservation in our Vlasov-Poisson solver and the method is convergent. We have verify that the energy remains unchanged to machine precision from the full resampling algorithm by printing out the second moment before and after resampling and use a very tight tolerance in the pseudoinverse solve.
6.1 Nonlinear Landau damping, $A=0.5$
The nonlinear Landau damping test case, $A=0.5$, has been studied extensively in the literature (Kraus Table 5.1 tabulates several results of previous work [19]). Cheng and Knorr test with a $32\times 128$ cell continuum grid solver and a time step of $\frac{1}{8}$ that is not only quantitatively similar to our results (Figure 3(d)) but qualitatively similar (Figure 4 in [7] and Figure 5 in [21] shows $\gamma_{1}$ measured with the first to peaks, and the fourth peak while the third is low). Figure 3 show convergence studies on the amplitude of the electric field for time step, resampling rate and particle and continuum grid resolution.
@@
These convergence tests establish parameters for an highly resolved analysis in Figure 3(d) that agrees well with Cheng and Knorr: $\gamma_{1}=-0.281$ as compared to our rate of $-0.278$, and a growth rate of $0.084$ vs our rate of $0.091$. Figure 3(b) show convergence with time step. The differences are only visible in the rebound stage and approximately second order accuracy is observed. The modest effect of resampling in this nonlinear case is observed in Figure 3(c), where resampling appears to increase the amplitude in the rebound stage. In comparison with the (clean) continuum results of Kraus (Figure 5.11 [19]) the onset of rebound is about the same but slope of the rebound stage with our code, without resampling appears to be higher and this slope is $\gamma_{2}=0.07458$ in Kraus and $\gamma_{2}=0.07885$ in Figure 3(c).
@@ 6.2 Linear Landau damping
A linear case of Landau damping, $A=0.0001$, demonstrates the potential of the projection algorithm. Figure 4 show the electric field amplitude with a variety of resampling rates with $Q2$ spaces in both space and velocity.
Figure 4: $E_{max},A=0.0001$, long time simulation with and without resampling, $Q2-Q2$ spaces
@@
No resampling is clearly very noisy and all of the resampling tests are free of noise, demonstrating the potential of this projection method for quiet, long time simulations. Resampling does seem to kick the plasma into growth phase faster, given that tests with a high resampling rate start to grow earlier. The damping rate of $\gamma_{1}=-0.15348$ agrees with theory, $\gamma_{1}=-0.153$, reported in [19].
@@ 7 Conclusion
The paper develops a new approach to particle resampling that uses a conservative projection, a pseudoinverse, to map any distribution of particles to essentially any other particle grid while conserving all moments up to the degree of polynomial that the projection function space can represent exactly. This method is evaluated with a simple projection grid, the original grid, on standard Landau damping problems. Tests where noise is problematic, the linear cases, show that resampling reduces noise considerably and coherent dynamics are maintained for long times, where as the solution becomes essentially all noise without resampling.
There area several areas of future work in developing adaptivity strategies for this method, for both particles and the projection grid, such as limiting the numerical entropy inherently generated in resampling, using adaptive continuum grid adaptivity (AMR) and using a modified original grid with particle splitting and coalescing techniques.
@@
With entropy measures, from our particle Landau collision operator [33], we can determine the continuum grids required for resampling or a continuum collision operator [3] to keep entropy generation by the projection well below the entropy generated in the collision operator.
AMR
AMR continuum grids in velocity space are available in PETSc [3], and regular particle grids on each cell of an adapted grid, like a cubed sphere, is a path for generating adapted particle grids through continuum mesh adaptation as is done in Myers et al. [31].
@@
These ideas, developed in many particle resampling methods, would allow for an incremental modification of the particle mesh to minimize cost and perhaps impact the physics less. We do observe that resampling increases the onset of rebound in Figure 4, which may imply that the physics is disturbed by the resampling and should therefor be minimized.
@@
Understanding the effects of resampling on physics, beyond conserving moments and other structure like entropy stability, requires experimentation, verification and validation, with more complex models like the Ion Temperature Gradient (ITG) instability.

### v5

@@ 4.2 Particle grids and finite element order
The phase space continuum grids in this work are on regular 90 degree lattice with an option for simple r-refinement in velocity space. Cartesian sub-particle grids are defined in each phase space cell, similar to Lapenta (Figure 1, [27]). For simplicity, the original grid is used for resampling in this work and adaptivity strategies are left for future work. As discussed in §3.1, with periodic boundary conditions in the spatial dimension and natural boundary conditions in velocity dimension we use at least $Q+1$ particles in each dimension in each phase space cell for stability of the pseudoinverse.
@@
Data with r-refinement grids have “graded” in the title of the plot and uniform grid data have “uniform” in the title. Similarly, the order of the finite element space in the data is encoded in the title, for example, $Q_{x}1-Q_{v}2$ uses linear $Q1$ (or $P1$) elements in real space and quadratic elements in velocity space.
@@ 5 Finite element version of the direct remap method
Colella et al. developed a method that can be viewed as a version of our algorithm in a finite volume context. In this algorithm continuum space points define the new particle points directly allowing for a trivial map from the grid to particles [43, 32, 33]. That is, $M_{p}=I$ in (3) and the pseudoinverse vanishes. Cubic splines are used for particle deposition on Cartesian grids, demonstrate high-order convergence [32], and use phase-space adaptive mesh refinement (AMR) grids [33]. This method should conserve energy because the cubic spines can represent a second order polynomial exactly and the direct remap obviously conserves all moments.
The test harness supports a direct map method in a vertex-centered, finite element context. Linear $Q1$ elements are used in the velocity and space dimension on Cartesian grids. Particles are placed at vertices and the domain is doubly periodic (to make $M_{p}$ square), which is immaterial given that there should be negligible density at the velocity boundary. The salient feature of this construct is that the particle mass matrix is the identity, $\overline{M}_{p}=I$, and the pseudoinverse vanishes. These experiments are intended to observe the qualitative effect of high-order grid to particle mapping. Quantitatively, the direct map approach conserve energy to about four digits with $Q1$ in velocity space (see the src/A.01 directory in the repository). These $Q1$ elements in velocity space with two particles per cell should, in theory conserve momentum only.
Figure 2: Particle electric field amplitude linear Landau damping, $A=0.01$, no resampling (blue and noisy), a direct remapping finite element version of Myers et al. (magenta), and the projection method (red)
Direct map resampling is tested with a linear Landau damping, $A=0.01$, problem with a $128\times 256$ particle grid and $V_{max}=6.0$ and results in Figure 2 agree well with amplitude reported in [32] (Figure 3.1). $Q_{1}-Q_{1}$ is used for the projection method with one particle per cell, or vertex, to mimic the direct map method, and the projection method uses a $2\times 2$ particle grid per cell and half as many cells in each dimension to maintain the same number of particles for all three tests and to ensure stability of the pseudo-inverse. This data shows the efficacy of resampling in that it suppresses the noise observed without resampling for the $A=0.01$ case that was also observed by Colella et al. and some increase in stability is observed with the pseudoinverse technique.
@@ 6 Numerical experiments with pseudoinverse resampling
This section investigates this projection resampling method on a linear and a nonlinear Landau damping tests. One issue to be addressed in a PIC method is ensuring a $C^{0}$ electric field, which our $C^{0}$ discretization of the Poisson equation does not provide given that one order of continuity is lost in the gradient of the potential. One can use an $H(div)$ Poisson solver or $C^{1}$ finite elements, but with $C^{0}$ elements we project the field to the vertices and then back to the particles, resulting a $C^{0}$ electric field. This method helps to stabilize our method, is all but required for $Q2$ elements in space. These instabilities could potentially be addressed with a quite start method [38]. Mesh adaptivity in velocity space [3], may act as a type of quite start, and is a subject of future work.
6.1 Nonlinear Landau damping
The nonlinear Landau damping test case, defined as setting $A=0.5$, has been studied extensively in the literature (Kraus Table 5.1 tabulates several results of previous work [21]). Cheng and Knorr test with a $32\times 128$ cell continuum grid solver and a time step of $\frac{1}{8}$ that is not only quantitatively similar to our results (Figure 4(b)) but qualitatively similar (Figure 4 in [9] and Figure 5 in [23] shows $\gamma_{1}$ measured with the first two peaks, and the fourth peak while the third is low). Figure 3 and 4 show convergence studies on the amplitude of the electric field for time step, resampling rate and particle and continuum grid resolution.
@@ 6.1 Nonlinear Landau damping
(a) Convergence in resampling period
(b) Highly converged
Figure 4: Converge study of nonlinear Landau damping, $A=0.5$
These convergence tests establish parameters for an highly resolved analysis in Figure 4(b) that agrees well with Cheng and Knorr: $\gamma_{1}=-0.281$ as compared to our rate of $-0.278$, and a growth rate of $0.084$ vs our rate of $0.090$. Figure 3(b) shows convergence study in time step where convergence is observed in the rebound region. The modest effect of resampling in this nonlinear case is observed in Figure 4(a), where resampling appears to increase the amplitude in the rebound stage.
@@ 6.2 Linear Landau damping
The linear case of Landau damping, $A=0.0001$, is used here to demonstrate the potential of the projection algorithm. Figure 5 shows the electric field amplitude with a variety of resampling rates with $Q2$ spaces in both real space and velocity space. A damping rate of $\gamma_{1}=-0.15348$ is observed, which agrees with theory reported in [21].
Figure 5: $E_{max},A=0.0001$, resample rate, and without resampling, “convergence” test
@@
With no resampling, the results are clearly very noisy while all of the tests with resampling, regardless of the rate, are free of noise. Resampling does, however, seem to kick the plasma into the growth phase faster, given that tests with a high resampling rate start to grow earlier.
7 Refinement studies and long time stability
[…]
Figure 6 shows a refinement study in the number of particles per cell which does not effect the location of the rebound. We further note that the level of the rebound converges at about $1M$ particles as is shown in Figure 8(b) and Figure 6 box “(b)”. The amplitude of the second rebound ($T=100$ in units of inverse plasma frequency $[\omega_{p}^{-1}]$) is very close for all but the lowest particle count ($16K$) model and this alternating pattern slowly decays leaving coherent and noise-free waves (see detail at $T=500$ Figure 8(c)). However, the $16K$ particle model starts to heat up and continues to grow at longer times ($T=1,000$ data is available in the repository). The $65K$ and $259K$ models remain stable with clean oscillations up to $T=1,000$. The observed change in amplitude of the rebound may be explained by the plasma’s increased ability to absorb wave energy as the particle density is increased. We expect this to converge monotonically but the additional presence of statistical noise in the system, as well as the rebound’s apparent dependence on its state just before rebounding may offer an explanation as to why we do no observe a clean convergence structure.
Figure 6: $E_{max},A=0.0001$, Refinement in particles per cell long time simulations
Figure 7 shows refinement studies in grid density with constant number of particles per phase-space cell. An increase in cell count delays the onset of subgrid modes and suppresses the rebound with the largest case almost entirely suppressed before $T=500$, when a large amplitude subgrid mode erupts and the continues with high-frequency, yet coherent, oscillations.
Figure 7: $E_{max},A=0.0001$, Refinement in grid density long time simulations
(a) $\Delta t$ convergence ($\Delta t=0.01$ and $\Delta t=0.025$), $3.6M$ particles Fig.7
(b) Convergence detail Fig. 6 (box)
(c) Long time stability Fig. 6 (box)
Figure 8: $E_{max},A=0.0001$, long time simulations
The choice in timestep is based on previous works with PETSc-PIC [13, 14] and Figure 8(a) shows little difference with a smaller time step than used in these studies.
8 Side effects of resampling
Figure 4(a) demonstrates that the dynamics of the plasma is not highly sensitive the resampling period within a broad range (eg, $1-8$ $[\omega_{p}^{-1}]$), but that rebound starts a little earlier with higher resampling rates indicating that resampling is disturbing the plasma to some extent. To investigated this effect, Figure 9 (left) shows a linear example with resampling every time step and increasing resampling periods. The highest resampling rates clearly lead to instabilities and these instabilities demonstrate classic growing oscillations up to a small amplitude, see detail in Figure 9 (right), that is growing in mean value and plateaus at about $1.0$. Perhaps stability analysis could provide insight on this issue.
Figure 9: $E_{max},A=0.0001$, resampling period “anti-convergence” study (left), instability detail (right)
While this instability at high resampling rates is not a practical problem, per se, in that all of our tests are stabilized with much lower resampling frequencies, this data does indicate that resampling can have adverse effects. This test problem is highly idealized and does not provide convincing experimental evidence that the adaptivity strategy of simply mapping make to the original grid will be robust in practice. More sophisticated adaptivity strategies are the subject of future work and would require more scientifically relevant models such as problems in magnetic reconnection in at least $2X+3V$.
8.1 Discussion
Figure 7 sheds light onto the complex relationship and dependency between all the simulation variables, i.e. resampling rate, spatial and velocity grid density and time step. It is difficult to draw definitive conclusions about what exactly is driving the dynamics of the electric field damping and rebounding. As discussed in §7, the grid resolution plays a large role in reducing the interaction of subgrid modes with the physical waves, while the particle density reduces stochastic, noise-driven heating effects in the plasma. These effects do not, however, converge monotonically suggesting that they are linked in some way. This requires further study in future work.
The inclusion of collisions may also provide a smooth mechanism. The collision operator acts as a controlled, physically motivated diffusion term in the kinetic equations. This diffusive term continuously smooths out any build up of high-frequency components in the distribution function which can lead to spurious rebounds. In other algorithms (Fig. 5.5 [21]), clean decays to machine precision have been shown using collisions. Collisions must be used with some level of caution, however. Much of the essential collisionless interactions in the Landau damping test, in particular in the nonlinear regime, rely on fine-scale interactions between the particles and waves. Introducing a collision operator can have the effect of smoothing out those essential fine-scale interactions which, if too strong, will disrupt the fundamental physics of the system. Recent work in the PETSc library [2, 35, 1, 3] has provided a variety of collision algorithms that can be included in the PETSc-PIC framework in future work. Other methods, such as [47], have reported clean decay without collisions, but the details of the numerical method are not entirely clear and some form of dissipation my be present.
@@ 9 Conclusion
The paper develops a new approach to particle resampling that uses a conservative projection, a pseudoinverse, to map any distribution of particles to essentially any other distribution of particle while conserving all moments up to the degree of polynomial that the projection function space can represent exactly. This method is evaluated with a static particle and continuum Cartesian grids, simply remapping to the original particle grid, on standard Landau damping test problems. Tests where noise is problematic, the linear cases, show that resampling reduces noise considerably and coherent dynamics are maintained for long times, where as the solution becomes essentially all noise without resampling.
This work suggests several areas of future work, namely:
Adaptivity strategies
Developing strategies to optimize the particle grids with respect to minimizing disturbances to the physics from resampling.
@@
With entropy measures, from our particle Landau collision operator [35], we can determine the continuum grids required for resampling or a continuum collision operator [3, 35] to keep entropy generation by the projection well below the entropy generated by the physics.
Continuum grid AMR
AMR in velocity and regular particle grids on each cell of an adapted grid, like a cubed sphere [3], is a path for generating adapted particle grids through continuum mesh adaptation.
@@
These ideas, developed in many particle resampling methods, would allow for an incremental modification of the particle mesh to minimize cost and perhaps have less impact the dynamics.
@@
Understanding the effects of resampling on physics, beyond conserving moments and other structure like entropy stability, requires experimentation with more complex models such as the Ion Temperature Gradient (ITG) instability to understand the robustness of these ideas.